A ball is kicked from a location (7,0, -8) (on the ground) with initial velocity (-9, 14, -4) m/s. The balls speed is low enough that air resistance is negligible. (a) What is the velocity of the ball 0.2 seconds after being kicked? (Use the Momentum Principle!) (b) In this situation (constant force), which velocity will give the most accurate value for the location of the ball 0.2 seconds after it is kicked? the final velocity of the ball the initial velocity of the ball the arithmetic average of the initial and final velocities (c) What is the average velocity of the ball over this time interval? m/s (d) Use the average velocity to find the location of the ball 0.2 seconds after being kicked. Now consider a different time interval: the interval between the initial kick and the moment when the ball reaches its highest point. We want to find how long it takes for the ball to reach this point, and how high the ball goes. (e) What is the y-component of the balls velocity at the instant when the ball reaches its highest point (the end of this time interval)? m/s (r) Fill in the missing numbers in the equation below (update form of the Momentum Principle: mvyt myFnet, At) (9) How long does it take for the ball to reach its highest point? 쏘. (h) Knowing this time, first find the y-component of the average velocity during this time interval, then use it to find the maximum height attained by the ball. max
Answer #1

In this case only velocity along y will change

No change in velocity along x and z

velocity along y will be=14-gt

=14-10(.2)

=12m/s

(g=10m/s^2)

Velocity after 0.2s=(-9,12,-4)m/s

b)initial velocity of the ball

c)final position of the ball

initial position of the ball=7,0,-8

final position of the ball along x=7-9*.2

=7-1.8=5.2m

final position of the ball along y=14t-1/2gt^2

=14*.2-5*.2*.2

=2.8-.2

=2.6m

final position of the ball along z=-8-4*.2

=-8.8

final position(5.2,2.6,-8.8)

initial position(7,0,-8)

total displacement=3.26

Total time=.2s

Avg velocity=3.26/.2

=16.3m/s

d)final position(5.2,2.6,-8.8)